# Foundational Policy Acquisition via Multitask Learning for Motor Skill Generation

### v3

@@ Abstract
In this study, we propose a multitask reinforcement learning algorithm for foundational policy acquisition to generate novel motor skills. Inspired by human sensorimotor adaptation mechanisms, we aim to train encoder-decoder networks that can be commonly used to learn novel motor skills in a single movement category. To train the policy network, we develop the multitask reinforcement learning method, where the policy needs to cope with changes in goals or environments with different reward functions or physical parameters of the environment in dynamic movement generation tasks. Here, as a concrete task, we evaluated the proposed method with the ball heading task using a monopod robot model. The results showed that the proposed method could adapt to novel target positions or inexperienced ball restitution coefficients. Furthermore, we demonstrated that the acquired foundational policy network originally learned for heading motion, can be used to generate an entirely new overhead kicking skill.
@@
Robot learning; Meta reinforcement learning; Domain Adaptation; Policy selection
publicationid: pubid: $0$–$00$ 2024 IEEE
@@ I Introduction
Multitask reinforcement learning (multitask RL) has been gaining attention as a promising tool to develop human-level skillful agents. In particular, the meta-reinforcement learning (meta-RL) approach that updates shared policy parameters among different policies for different tasks has become a significant research topic for optimizing the parameters of the policies to cope with the multitask RL problems [1, 2, 3, 4]. In this study, we propose to adopt the idea of multitask RL to acquire a foundational motor skill that can adapt to generate novel motor skills by using encoder-decoder networks. As the computational neuroscience studies stated [5, 6, 7], human sensorimotor adaptation mechanisms may contain explicit motor skill learning processes and non-intentional implicit adaptation processes to circumstances. Inspired by these sensorimotor adaptation systems, we propose using a multitask RL framework to acquire a foundational policy that can take the role of implicit adaptation and optimize latent variables of the decoder network to generate novel skills as the explicit motor learning process. Here, we consider the acquisition of a football skill as the target task since football is one of the most popular sports in the world [5]. Furthermore, for example, the ball can fly in completely different directions with a similar heading motion by slightly changing the hitting position, the force applied to the ball, or when the friction of the ball is different. In such cases, continuous task changes may have infinite numbers of task goals, and preparing policy networks for all these tasks becomes impossible. The previous meta-RL framework evaluated the obvious task changes, e.g., the door opening and pick-and-place, and the implicit task change task is insufficient to evaluate and develop.
Therefore, this study proposes a training procedure for policies represented by deep neural networks to cope with the multitask RL problems with the task changes. We were particularly interested in task changes within the same movement category: heading a ball to different goal locations. As the concrete learning framework, we developed a novel skill acquisition method for the contextual Markov decision process (cMDP) to cope with the multitask problems based on a variational inference framework [3]. We demonstrated that selecting a base network model according to the learning performance with randomized task parameters is the key to successful adaptation to multiple reward and environmental settings. Furthermore, our proposed method simultaneously acquires the optimal policy for each context and encoder to embed context variables that correspond to the task settings. The context-related latent variable was estimated to generate skills to handle the novel tasks: unknown rewards and environmental settings. We showed that the Bayesian optimization method can be efficiently used to update the latent variable in the adaptation procedure in cases where the task-relevant variable space is low dimensional (see also Fig. 1).
(a) Multitask heading
(b) Throw in
(c) Hitting ball
Figure 1: Proposed three-phase multitask learning method. The proposed method comprises three phases: acquisition, selection, and generation. In the foundational policy acquisition phase, multiple candidate policies are generated under multiple domains. (a) The $i$-th candidate policy $\pi_{i}$ is trained under the $j$-th domain. In the policy selection phase, one policy is selected from the candidate policy set for the skill generation phase. Low-dimensional latent variables are optimized for adapting to novel tasks and environments in the skill generation phase. (b) The policy selection phase chooses a policy based on the statistical distance from candidate policies $\Pi$. (c) The latent variable $z$ is optimized based on the task performance of the new domain. Multitask heading problem. Monopod robot model tries to acquire policies to hit the ball back to multiple goals in multiple environments. (d, e, f) Implicit multitask heading. Start position, goal position, and the coefficient of the restitution are varied.
@@
We evaluated our proposed policy adaptation method using the monopod motor control task. In this task, as depicted in Fig. 1d-f, the monopod model tries to hit the ball back to the target point while maintaining the posture until it hits the ball. To show the adaptation performance of the proposed method, the target points, as well as the coefficient of restitution that defines the physical interaction between the ball and the monopod model, were changed. We demonstrated that the monopod model was successfully adapted to the implicitly changing task settings. The contributions of our study can be summarized as follows:
@@
  • We developed a policy adaptation method that can cope with implicit multitask problems. The proposed method is composed of three phases: 1) foundational policy acquisition, 2) policy selection, and 3) skill Generation, as shown in Fig. 1a-c.
@@
  • We demonstrated that selecting the base policy from the policy pool derived in the foundational policy acquisition phase in terms of learning performance is the key to successful adaptation.
@@
  • We showed that the monopod model was able to hit the ball back to novel target points with different physical interaction settings using the proposed skill generation method.
@@
The remainder of this paper is organized as follows: In section II, we described the related studies. In section III, we introduced the proposed adaptation method. In section IV, we explained the experimental settings for the monopod model used for evaluating the proposed method. In section V, we presented the results. In section VI, we discussed the comparison of the proposed method with the standard approach as well as the limitations of the proposed method. Finally, in Section VII, we presented the conclusions of the study.
@@ II Related Works
II-A Goal-conditioned RL
To acquire a policy that can cope with multiple task goals, goal-conditioned RL methods have been developed [8, 9, 10, 11, 12]. By acquiring a goal-conditioned value function, the learned policy can generate optimized behaviors according to the continuously changing task goals [9, 13], and [14]. Moreover, some studies proposed to use an image as a task goal [10, 13]. However, these approaches essentially just select a policy corresponding to a specified task goal and cannot handle implicitly changing task goals. This study proposes a three-phase learning and adaptation procedure for coping with implicitly changing environmental settings and task goals.
@@ II-B Data-driven System Identification
To cope with a novel environment by an online data-driven approach, adaptive control methods have been widely studied [15]. If we cannot rely on a parametric environmental model, the model is usually built based on domain-specific knowledge. In that case, we can use an approximated model, possibly represented by a neural network, to bridge the gap between the current estimated and the actual environmental settings. Our early studies proposed collecting data by first using a robust policy that can deal with the novel environment and further improving the policy based on the approximated physical model by using the newly acquired data from the novel environment [16, 17]. Recently, instead of using the robust policy, acquiring a policy network that takes physical parameters as inputs has been proposed [18]. The environment-aware policy was acquired under a variety of environmental settings in the training phase. In the test phase, the physical parameters were identified to improve the performance of the policy further. This approach has been further explored using a low-dimensional representation of the physical parameters. In the training phase, the low-dimensional latent parameter space was extracted. Then, in the test phase, the latent parameters were identified through interaction with the novel environment. The above approach was adopted to deal with the simulation-to-real gap [19, 4]. In this study, we applied these system identification approaches to the implicit multitask RL problem. We showed that the policies can be adapted not only to novel environments but also to novel tasks.
II-C Policy Selection
In this study, we proposed an approach to find an adaptable base policy for unknown tasks and environments. We must consider the policy selection problem [20] to detect the adaptable policy. Our proposed method selected the adaptable base policy by evaluating the learning performances of each policy generated in the foundational policy acquisition phase.
II-D Derivative-free Optimization
@@ III Method
In this section, we first defined the cMDP as multitask reinforcement learning, in which the task goal or environments change implicitly. We then introduced the novel variational inference-based reinforcement learning method to cope with the cMDP.
@@ III-A Solving MDP through KL-divergence minimization
$\max_{\pi}\ \mathbb{E}_{\tau|\pi}\left[\sum_{t=0}^{T}\left(r(s_{t},a_{t})-\alpha\log\pi(a_{t}|s_{t})\right)\right],$ (1)
@@ III-B Solving cMDP
This subsection introduced the proposed variational inference-based RL method for solving the contextual MDP. A cMDP $\{\mathcal{A},\mathcal{S},\mathcal{C},r,p,p_{0}\}$ has the state space $\mathcal{S}$, action space $\mathcal{A}$, context space $\mathcal{C}$, reward function $r(s,a,c)\in\mathbb{R}$, state transition probability $p(s^{\prime}|s,a,c)\in\mathbb{R}$, and initial distribution of the state and context $p_{0}=p(s_{0})p(c)\in\mathbb{R}$. State transition probability and reward functions are conditioned with a vector $c$ in context space $\mathcal{C}$. To acquire the optimal policy in cMDPs, we trained a policy $\pi(a|s,z)$, which depended on states $s$ and the latent variable $z$. Here, the latent variable $z$ is a context-relevant feature vector sampled from an encoder distribution $q(z|c)$. In other words, the encoder converts the current context variables $c$ to the latent variable $z$.
@@
The variational distribution $q$ is the probability of trajectory $\tau$ under context $c\sim p(c)$ and latent variable $z\sim q(z|c)$. Likewise, a generative model $p$ is the target distribution close to the variational distribution $q$; $\rho(z)$ is the prior distribution; and $\alpha$ is the scale parameter. The variational distribution and generative model share the context distribution $p(c)$ and transition distribution $p(s_{t+1}|s_{t},a_{t},c)$. Therefore, cumulative rewards with entropy regularization by canceling out shared distribution were obtained as follows:
$\alpha\mathrm{KL}[q|p]$ $=\alpha\mathbb{E}_{q}[\log q-\log p]$
@@ III-C Learning Algorithm
The proposed multitask learning algorithm is composed of three phases: generation, selection, and adaptation phases, as shown in Fig. 1.
@@ III-C1 Foundation Policy Acquisition
We then updated the policy and the encoder using the KL divergence introduced in (10) as follows:
@@
$\phi$, $\psi$, and $\eta$ represent the parameters of the Q function, the policy, and the encoder, respectively. Fig. 2 shows the architecture of each network.
The prior of the latent variable $\rho(z)$ was derived by a standard multivariate normal distribution. The replay buffer stored the tuple $\{s,a,r,s^{\prime},c\}$.
@@
19:  Output $\phi_{\{1,2\}},\psi,\eta$
@@ III-C2 Selection
One of the key ideas of our proposed method was to select a policy model based on the return values in the foundational policy acquisition phase. As a well-known issue, the task performance of a policy highly depends on an exploration procedure determined by a random seed. Thus, in this paper, we proposed the generation of multiple policy models in the offline RL manner to find a high-quality policy that can adapt to a novel task from the foundational policy models. The detailed selection procedure is explained in Algorithm 2. The average reward over the $k$-th episode, $l_{k}$, was stored in the generated $k$-th policy model. We then selected the policy model based on the KL loss in Eq. (10) by substituting the value function with Eq. (7).
@@
2:  Setup randomized domain with $\mathrm{nun}_{c}$ contexts $\{c\}$
@@
4:     Train $\pi_{k}$ with contexts $\{c\}$triangleright Generation Phase
5:     Rollout $\pi_{k}$ and collect $R(c):=\sum_{t}{\gamma^{t}(r_{t}-\alpha\log\pi_{t})}$
6:     $l_{k}=\frac{1}{\mathrm{num}_{c}}\sum_{c}\left[\beta\log\frac{q(z|c)}{\rho(z)}-R(c)\right]$ in Eq. (10)
@@ III-C3 Skill Generation
The encoder network learns the latent variable space, where the context variables are embedded in this latent space. In the skill-generation phase, we only optimized the latent variable such that the policy could cope with a novel domain.
@@
Skill-generation encoder $P(*|z)$ is represented with multivariate normalized distribution $\mathcal{N}(z,\mathbf{I})$. As we adopted a one-dimensional latent space, black-box optimization methods were utilized. As a concrete optimization method, we chose the TPE, which has been used as a standard tool for searching hyperparameters [23], because it shows empirically better optimization performance than other types of Bayesian optimization methods, such as the GPs and random search methods [23, 24]. To solve the optimization problem, TPE estimates the conditional distribution $p(z|J)$ from sampled data $\mathcal{D}=\{(z,J)\}$. We adopted the off-the-shelf implementation provided in the hyperparameter optimization framework, Optuna [27]. The mean $z$ of encoder distribution is suggested by TPE for each iteration, as shown in Algorithm 3. The algorithm rollouts the policy based on the suggested latent variable $z_{k}$ to collect the return value $J(z_{k})$ and stores it in the replay buffer. The TPE updates the posterior distribution model for the next suggestion. The rollout will be iterated $k_{\mathrm{max}}$ times. Finally, the algorithm returns the best value $z^{*}$ stored in the replay buffer $\mathcal{D}$.
@@
1:  $\pi\leftarrow$ trained policy in Policy Search phase
@@
6:     $J(z_{k})\leftarrow$ Rollout policy $\pi(a|s,z_{k})$ in Eq. (14)

### v4

@@ Abstract
In this study, we propose a multitask reinforcement learning algorithm for foundational policy acquisition to generate novel motor skills. Learning the rich representation of the multitask policy is a challenge in dynamic movement generation tasks because the policy needs to cope with changes in goals or environments with different reward functions or physical parameters. Inspired by human sensorimotor adaptation mechanisms, we developed the learning pipeline to construct the encoder-decoder networks and network selection to facilitate foundational policy acquisition under multiple situations. First, we compared the proposed method with previous multitask reinforcement learning methods in the standard multi-locomotion tasks. The results showed that the proposed approach outperformed the baseline methods. Then, we applied the proposed method to the ball heading task using a monopod robot model to evaluate skill generation performance. The results showed that the proposed method was able to adapt to novel target positions or inexperienced ball restitution coefficients but to acquire a foundational policy network, originally learned for heading motion, which can generate an entirely new overhead kicking skill.
@@
Robot learning; Multitask reinforcement learning; Domain adaptation; Policy selection
@@ I Introduction
Multitask reinforcement learning (multitask RL) has been gaining attention as a promising tool to develop human-level skillful agents. In particular, the meta-reinforcement learning (meta-RL) approach that updates shared policy parameters among different policies for different tasks has become an important research topic for optimizing the parameters of the policies to cope with the multitask RL problems [1, 2]. Brain-inspired frameworks for multitask learning problems have also been explored [3, 4]. In this study, we propose adopting the idea of multitask RL to acquire a foundational motor skill that can generate novel motor skills by using encoder-decoder networks. As the computational neuroscience studies stated [5], [6], [7], human sensorimotor adaptation mechanisms may contain explicit motor skill learning processes and non-intentional implicit adaptation processes to circumstances. Inspired by these sensorimotor adaptation systems, we propose using a multitask RL framework to acquire a foundational policy that can take the role of implicit adaptation and optimize latent variables of the decoder network to generate novel skills as explicit motor learning.
First, to evaluate the multitask learning performance, we compared the proposed method with previous meta-RL methods in the standard multi-locomotion benchmarks [8, 2]. Then, we applied the proposed method to the ball heading task using a monopod robot model to evaluate skill generation performance since football is one of the most popular sports in the world [5]. Furthermore, for example, the ball can fly in completely different directions with a similar heading motion by slightly changing the hitting position, the force applied to the ball, or when the friction of the ball is different. In such cases, continuous task changes may have infinite numbers of task goals, and preparing policy networks for all these tasks becomes impossible.
As illustrated in Fig. 1, this study proposes a three-phase training procedure for policies represented by deep neural networks to cope with the multitask RL problems with the task changes. Unlike the previous meta-RL methods, our proposed method can generate a novel movement skill based on a newly developed variational inference framework to train the policy network models.
Figure 1: Proposed three-phase multitask learning method. Proposed pipeline comprises three phases: acquisition, selection, and generation. In the foundational policy acquisition phase, multiple candidate policies are generated under multiple tasks. (a) $i$-th candidate policy $\pi_{i}$ is trained under multiple tasks. (b) Policy selection phase chooses a policy from the policy set $\Pi$ based on the performance index defined in Eq. (15). (c) Context-related latent variable $z$ was estimated to generate skills to handle the novel tasks: unknown rewards and environmental settings. We showed that the Bayesian optimization method can be efficiently used to update the latent variable in the skill generation.
(a) Target direction
(b) Target veloctiy
(c) Multitask heading
(d) Throw in
(e) Hitting ball
Figure 2: Multitask environments to generate a variety of dynamic movements. (a) Half-Cheetah-Dir domain. Goal is to move forward or backward as fast as possible. (b) Half-Cheetah-Vel domain. Goal is to reach a target velocity [8, 2]. (c, d, e) Implicit multitask heading. Start position, goal position, and the coefficient of the restitution are varied.
@@
We first evaluated our proposed approach in the standard multi-locomotion benchmarks illustrated in Figs. 2(a), (b). In the Half-Cheetah-Dir domain, the goal is to move forward or backward as fast as possible with using the Half-Cheetah model. In the Half-Cheeetah-Vel domain, the goal is to reach a target velocity [8, 2]. We then evaluated the skill generation performance of our proposed policy adaptation method using the monopod motor control task. In this task, as depicted in Figs. 2(c), (d), (e), the monopod model tries to hit the ball back to the target point while maintaining the posture until it hits the ball. To show the skill generation performance of the proposed method, the target points, as well as the coefficient of restitution that defines the physical interaction between the ball and the monopod model, were changed. We demonstrated that the monopod model was successfully adapted to the implicitly changing task settings. Furthermore, we showed that the new movement skill could be generated by using the acquired foundational model. The contributions of our study can be summarized as follows:
@@
  • We developed a policy adaptation pipeline that can cope with implicit multitask problems. The proposed method is composed of three phases: 1) foundational policy acquisition, 2) policy selection, and 3) skill generation, as shown in Fig. 1.
@@
  • We demonstrated that selecting the base policy from the policy pool derived in the foundational policy acquisition phase in terms of learning performance is the key to successful skill generation. To evaluate the multitask learning performance, we compared our proposed method with previous meta-RL methods in the standard meta-locomotion domains. The results showed that the proposed approach outperformed the baseline methods.
@@
  • We showed that the monopod model was able to hit the ball back to novel target points with different physical interaction settings using the proposed skill generation method. We further demonstrated that the monopod model was able to generate a novel kicking skill that was not explicitly acquired in the foundational policy acquisition phase.
@@
The remainder of this paper is organized as follows: In section II, we described the related studies. In section III, we introduced the proposed adaptation method. In section IV, we explained the experimental settings for the Half-Cheetah and monopod model used for evaluating the proposed method. In section V, we presented the results. In section VI, we discussed the comparison of the proposed method with the standard approach as well as the limitations of the proposed method. Finally, in Section VII, we presented the conclusions of the study.
@@ II Related Works
II-A Policy Learning Methods for Multiple Task Goals
To acquire a policy that can cope with multiple task goals, goal-conditioned RL methods have been developed [9],[10],[11],[12], [13]. Goal-conditioned approaches have also been explored in imitation learning frameworks [14, 15]. By acquiring a goal-conditioned value function, the learned policy can generate optimized behaviors according to the continuously changing task goals [10, 16], [17]. Moreover, some studies proposed to use an image as a task goal [11, 16]. However, these approaches essentially select a policy corresponding to a specified task goal and cannot handle implicitly changing task goals. This study proposes a three-phase learning and adaptation procedure for coping with implicitly changing environmental settings and task goals. As a part of the proposed pipeline, we considered policy selection problem [18] to derive a foundation policy. Concretely, the policy selection phase chooses a policy from a policy set based on the performance index defined by the Kullback–Leibler (KL) divergence between trajectory distributions (see Section III-B for details).
@@ II-B Data-driven System Identification
The online data-driven approach is used to cope with a novel environment, and adaptive control methods have been widely studied [19]. If we cannot rely on a parametric environmental model, the model is usually built based on domain-specific knowledge. In that case, we can use an approximated model, possibly represented by a neural network, to bridge the gap between the current estimated and the actual environmental settings. Our early studies proposed collecting data by first using a robust policy that can deal with the novel environment and further improving the policy based on the approximated physical model by using the newly acquired data from the novel environment [20, 21]. Recently, instead of using the robust policy, acquiring a policy network that takes physical parameters as inputs has been proposed [22]. The environment-aware policy was acquired under various environmental settings in the training phase. In the test phase, the physical parameters were identified to improve the policy performance. This approach has been further explored using a low-dimensional representation of the physical parameters. In the training phase, the low-dimensional latent parameter space was extracted. Then, in the test phase, the latent parameters were identified through interaction with the novel environment. The above approach was adopted to deal with the simulation-to-real gap [23, 24]. In this study, we applied these system identification approaches to the implicit multitask RL problem. We showed that the policies can be adapted not only to novel environments but also to novel tasks.
II-C Derivative-free Optimization
@@ III Method
In this section, we first defined contextual MDP (cMDP) as multitask reinforcement learning, in which the task goal or environments change implicitly. We then introduced the novel variational inference-based reinforcement learning method to cope with cMDP.
@@ III-A Solving MDP through KL-divergence Minimization
$\max_{\pi}\ \mathbb{E}_{\tau|\pi}\left[\sum_{t=0}^{T}r(s_{t},a_{t})-\alpha\log\pi(a_{t}|s_{t})\right],$ (1)
@@ III-B Solving cMDP
This subsection introduced the proposed variational inference-based RL method for solving cMDP. Contextual MDP $\{\mathcal{A},\mathcal{S},\mathcal{C},r,p,p_{0}\}$ has the state space $\mathcal{S}$, action space $\mathcal{A}$, context space $\mathcal{C}$, reward function $r(s,a,c)\in\mathbb{R}$, state transition probability $p(s^{\prime}|s,a,c)\in\mathbb{R}$, and initial distribution of the state and context $p_{0}=p(s_{0})p(c)\in\mathbb{R}$. State transition probability and reward functions are conditioned with a vector $c$ in a context space $\mathcal{C}$. To acquire the optimal policy in cMDPs, we trained a policy $\pi(a|s,z)$, which depended on states $s$ and the latent variable $z$. Here, the latent variable $z$ is a context-relevant feature vector sampled from an encoder distribution $q(z|c)$. In other words, the encoder converts the current context variables $c$ to the latent variable $z$.
@@
The variational distribution $q$ is the probability of trajectory $\tau$ given context $c\sim p(c)$ and latent variable $z\sim q(z|c)$. Likewise, a generative model $p$ is the target distribution close to the variational distribution $q$, and $\rho(z)$ is the prior distribution. The variational distribution and generative model share the context distribution $p(c)$ and transition distribution $p(s_{t+1}|s_{t},a_{t},c)$. Therefore, cumulative rewards with entropy regularization by canceling out shared distribution were obtained as follows:
$\alpha\mathrm{KL}[q|p]$ $=\alpha\mathbb{E}_{q}[\log q-\log p]$ (6)
@@ III-C Learning Algorithm
The proposed multitask learning algorithm is composed of three phases: acquisition, selection, and generation phases, as shown in Figs. 1 and 3.
(a) Foundational policy acquisition
(b) Policy selection
Figure 3: Foundational policy acquisition and policy selection. We demonstrated that selecting a foundational policy model according to the learning performance with randomized task parameters is the key to successful adaptation to multiple reward and environmental settings. Furthermore, our proposed method simultaneously acquires the optimal policy for each context and encoder to embed context variables that correspond to the task settings.
@@ III-C1 Foundation Policy Acquisition
We then updated the policy and the encoder using the KL divergence introduced in Eq. (11) as follows:
@@
Here, $\phi$, $\psi$, and $\eta$ represent the parameters of the Q function, the policy, and the encoder, respectively. Fig. 4 shows the architecture of each network. The prior of the latent variable $\rho(z)$ was derived by a standard multivariate normal distribution. The replay buffer $\mathcal{D}$ stored the tuple $\{s,a,r,s^{\prime},c\}$.
@@
19:  Output policy and encoder network weights $\psi,\eta$
@@ III-C2 Selection
One key idea of our proposed method was to select a policy model based on the return values in the foundational policy acquisition phase. As a well-known issue, the task performance of a policy highly depends on an exploration procedure determined by a random seed. Thus, in this paper, we proposed the acquisition of multiple policy models in the offline RL manner to find a high-quality policy that can adapt to a novel task from the foundational policy models. The detailed selection procedure is explained in Algorithm 2. The $k$-th candidate policy collects the sampled return value $R(c)$ for each context from a context set. The performance index $l_{k}$ defined in Eq. (11) is stored with the candidate policy $\pi_{k}$ for policy selection as follows:
$l_{k}$ $=\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}\left[\beta\log\frac{q(z|c)}{\rho(z)}+R(c)\right],$ (15)
$R(c)$ $=\sum_{t}\gamma^{t}\left(r_{t}-\alpha\log\pi_{t}(z)\right),z\sim q(z|c).$
We then select the policy with the highest performance index.
@@
2:  Sample context variables to define context set $\mathcal{C}$
@@
4:     Train $\{\pi_{k},q(z|c)\}$triangleright Foundational Policy Acquisition
5:     Rollout $\{\pi_{k},q(z|c)\}$ and collect $R(c):=\sum_{t}{\gamma^{t}(r_{t}-\alpha\log\pi_{t})}$ for each $c\in\mathcal{C}$
6:     Evaluated policy selection index $l_{k}$ in Eq. (15)
@@ III-C3 Skill Generation
The encoder network learns the latent variable space, where the context variables are embedded in this latent space. In the skill-generation phase, we only optimized the latent variable so that the policy could cope with a novel domain.
@@
Skill-generation encoder $P(*|z)$ is represented with multivariate normalized distribution $\mathcal{N}(z,\mathbf{I})$. We adopted black-box optimization methods to explore a one-dimensional latent space. As a concrete optimization method, we chose the TPE, which has been used as a standard tool for searching hyperparameters [27], because it shows empirically better optimization performance than other types of Bayesian optimization methods, such as the GPs and random search methods [27, 28]. To solve the optimization problem, TPE estimates the conditional distribution $p(z|J)$ from sampled data $\mathcal{D}=\{(z,J)\}$. We adopted the off-the-shelf implementation in the hyperparameter optimization framework, Optuna [31]. The mean $z$ of encoder distribution is suggested by TPE for each iteration, as shown in Algorithm 3. The algorithm collects rewards series from the deterministic action $a=\mu(s,z_{k})$, which is the pre-layer of noise injection in Fig. 4 based on the suggested latent variable $z_{k}$ to estimate the return value $J(z_{k})$ and store it in the replay buffer. The TPE updates the posterior distribution model for the next suggestion. The rollout will be iterated $k_{\mathrm{max}}$ times. Finally, the algorithm returns the best value $z^{*}$ stored in the replay buffer $\mathcal{D}$.
@@
1:  $\pi\leftarrow$ selected policy in Selection Phase
@@
6:     $J(z_{k})\leftarrow$ Rollout policy $a=\mu(s,z_{k})$ in Eq. (16)
@@
III-C4 Network models
The policy network was implemented using the multilayer neural network model. The network model had two hidden layers, shown in Fig. 4. The input variables to the network included the robot state $s$ and the latent variable of the encoder $z$. The encoder network embedded the context variable $c$ to the latent variable $z$ and implemented a neural network model with two hidden layers. In the foundational policy acquisition phase, the encoder and Q-function networks were trained using actual context variables $c$. Subsequently, in the skill generation phase, the encoder network was not used. The latent variable $z$ was estimated through adaptation trials.